OpenStudy (anonymous):

Sherry missed the lesson on normal distribution and needs to do her homework. Explain to Sherry how to use the mean and standard deviation of a normal distribution to determine the top 7% of the population.

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@jagr2713 @Librarian

OpenStudy (kirbykirby):

Well if you know the top 7%, you also know the lowest 93% of your distribution. So, if \(X\) is your normal distribution, you would be interested in the value (call it \(c\)) such that the values below \(c\) represent 93%, i.e. a probability of 0.93. So, \(P(X \le c)=0.93\) (You might have seen this also written as \(\Phi(c)=0.93\))

OpenStudy (kirbykirby):

But if you have a normal distribution, you can always transform it into a standard normal distribution by using the formula: \[Z=\frac{X-\mu}{\sigma} \] where \(\mu\) is the mean and \(\sigma\) is the standard deviation. Thus giving: \[P(X \le c)=P\left( \frac{X-\mu}{\sigma} \le \frac{c-\mu}{\sigma}\right)=P\left(Z\le \frac{c-\mu}{\sigma}\right) =0.93\]

OpenStudy (kirbykirby):

The formula for \(Z\) is called the z-score (or z-value). What you can do is look at the standard normal table for the cumulative probability value 0.93 and establish what z-score that value that occurs at. Then you'd equate the z-score with \(\dfrac{c-\mu}{\sigma}\) and solve for \(c\). Then you know that any score ABOVE \(c\) will represent the top 7% of the population